# Fourier integral operators on Orlicz modulation spaces

### v1

@@ Abstract.
We establish continuity and Schatten-von Neumann properties for Fourier integral operators with amplitudes in Orlicz modulation spaces, when acting on other Orlicz modulation spaces themselves. The phase functions are non smooth and admit second order derivatives in suitable classes of modulation spaces.
@@ 0. Introduction
The aim of the paper is to investigate continuity properties for Fourier integral operators with non-smooth amplitudes (or symbols), when acting on Orlicz modulation spaces. In particular we extend continuity properties of the Fourier integral operators in [7, 8, 58], where related results were obtained for such operators when acting on classical modulation spaces (of Lebesgue types). These earlier results are in turn extensions and generalizations of pioneering results and investigations performed by Boulkhemair in [5], where $L^{2}$-continuity of subclasses of our Fourier integral operators were established.
@@
is continuous. We notice that we may choose $\omega$ such that $M^{\infty,1}_{(\omega)}$ contain $S^{r}_{0,0}$. In particular, as announced above, we put less restrictions on the amplitudes compared to the assumptions above adapted to hyperbolic problems.
@@
Secondly, in (0.3) we allow $\omega$, $\omega_{j}$ and $v$ to belong to the large class $\mathscr{P}_{s}$ of weights which are moderate by subexponential functions of degree $s^{-1}<1$, while in [58] it is required that the involed weights should belong to the smaller class $\mathscr{P}$ of polynomially moderate weights. By imposing these two significant restrictions, then our Theorem 2.2 in Section 2, as well as (0.3), essentially takes the form [58, Theorem 2.2].
@@ 1. Preliminaries
In the section we recall some basic facts on Gelfand-Shilov spaces, Orlicz spaces, Orlicz modulation spaces, pseudo-differential operators and Wigner distributions. We also give some examples on Young functions, Orlicz spaces and Orlicz modulation spaces. (See Examples 1.9 and 1.15.) Notice that Young functions are fundamental in the definition of Orlicz spaces and Orlicz modulation spaces).
@@ 1.1. Gelfand-Shilov spaces
By [56, Theorem 2.3] it follows that the definition of the map $(f,\phi)\mapsto V_{\phi}f$ from $\mathscr{S}(\mathbf{R}^{d})\times\mathscr{S}(\mathbf{R}^{d})$ to $\mathscr{S}(\mathbf{R}^{2d})$ is uniquely extendable to a continuous map from $\mathcal{S}_{s}^{\prime}(\mathbf{R}^{d})\times\mathcal{S}_{s}^{\prime}(\mathbf{R}^{d})$ to $\mathcal{S}_{s}^{\prime}(\mathbf{R}^{2d})$, and restricts to a continuous map from $\mathcal{S}_{s}(\mathbf{R}^{d})\times\mathcal{S}_{s}(\mathbf{R}^{d})$ to $\mathcal{S}_{s}(\mathbf{R}^{2d})$. The same conclusion holds with $\Sigma_{s}$ in place of $\mathcal{S}_{s}$, at each occurrence.
In the following proposition we give characterizations of Gelfand-Shilov spaces and their distribution spaces in terms of estimates of the short-time Fourier transform. We omit the proof since the first part follows from [31, Theorem 2.7] and the second part from [56, Theorem 2.5]. See also [15] for related results. Here and in what follows, the notation $A(\theta)\lesssim B(\theta)$, $\theta\in\Omega$, means that there is a constant $c>0$ such that $A(\theta)\leq cB(\theta)$ holds for all $\theta\in\Omega$. We also set $A(\theta)\asymp B(\theta)$ when $A(\theta)\lesssim B(\theta)\lesssim A(\theta)$.
@@ 1.3. Orlicz Spaces
We refer to [60, Lemma 1.18] for the proof of the following lemma.
@@ 1.4. Orlicz modulation spaces
Next we explain some basic properties of Orlicz modulation spaces. The following proposition shows that Orlicz modulation spaces are completely determined by the behavior of the Young functions near origin. We refer to [61, Proposition 5.11] for the proof.
@@ 1.5. Kernel operators and Schatten-von Neumann classes
Throughout the paper, all Hilbert spaces are assumed to be separable, and observe that this is always the case for Hilbert spaces which are continuously embedded in $\mathscr{S}^{\prime}(\mathbf{R}^{d})$, in view of [42, Proposition 1.2]. However, we note that most parts of what is described here also hold when $\mathcal{H}_{1}$ and $\mathcal{H}_{2}$ are allowed to be non-separable.
@@ 2.2. Continuity for Fourier integral operators, when acting on modulation spaces
We usually assume that the involved weight functions satisfy
@@
Let $s>1$, $d_{1}=d_{2}=d$, $\Phi$ be a Young function such that $1<q_{\Phi}\leq p_{\Phi}<\infty$, $\varphi\in C(\mathbf{R}^{2d+m})$, $\omega,v\in\mathscr{P}_{s}(\mathbf{R}^{2(2d+m)})$ and $\omega_{1},\omega_{2}\in\mathscr{P}_{s}(\mathbf{R}^{2d})$ fulfill (0.2), (2.2) and (2.3), for some $d\hskip-0.80002pt\bar{}\hskip 1.99997pt>0$. Then the following is true:
  1. (1)
    the map $a\mapsto\operatorname{Op}_{\varphi}(a)$ from $\Sigma_{1}(\mathbf{R}^{2d+m})$ to $\mathcal{L}(\Sigma_{1}(\mathbf{R}^{d}),\Sigma_{1}^{\prime}(\mathbf{R}^{d}))$ extends uniquely to a continuous map from $M^{\infty,1}_{(\omega)}(\mathbf{R}^{2d+m})$ to $\mathcal{L}(\Sigma_{1}(\mathbf{R}^{d}),\Sigma_{1}^{\prime}(\mathbf{R}^{d}))$;
  2. (2)
    if $a\in M^{\infty,1}_{(\omega)}(\mathbf{R}^{2d+m})$, then the map $\operatorname{Op}_{\varphi}(a)$ from $\Sigma_{1}(\mathbf{R}^{d})$ to $\Sigma_{1}^{\prime}(\mathbf{R}^{d})$ extends uniquely to a continuous operator from $M^{\Phi}_{(\omega_{1})}(\mathbf{R}^{d})$ to $M^{\Phi}_{(\omega_{2})}(\mathbf{R}^{d})$. Moreover, for some constant $C$ it holds
@@
In the next result we suppose that $\|a\|$ is finite, where the norm $\|\,\cdot\,\|$ is defined, either by
@@
Then [58, Lemma 2.2] shows that
@@
The assertion (1) follows from assertion (1) in [58, Theorem 2.1].
We shall follow the proof of [58, Theorem 2.1] when proving (2). If $\Phi(t_{0})=0$ for some $t_{0}>0$, then $M^{\Phi}_{(\omega_{j})}(\mathbf{R}^{d})=M^{\infty}_{(\omega_{j})}(\mathbf{R}^{d})$, and the result follows from [58, Theorem 2.1]. Therefore suppose that $\Phi(t)>0$ when $t>0$. Then $\mathscr{S}(\mathbf{R}^{d})$ is dense in $M^{\Phi}_{(\omega_{j})}(\mathbf{R}^{d})$.
@@
The result follows from [58, Theorem 2.7] when $p_{\Phi}=q_{\Phi}$. Therefore we assume that $1<q_{\Phi}\leq p_{\Phi}<\infty$.

### v2

@@ Abstract.
We establish continuity, compactness and Schatten-von Neumann properties for Fourier integral operators with amplitudes in Orlicz modulation spaces, when acting on other Orlicz modulation spaces themselves. The phase functions are non smooth and admit second order derivatives in suitable classes of modulation spaces.
@@ 0. Introduction
The aim of the paper is to investigate continuity and compactness properties for Fourier integral operators with non-smooth amplitudes (or symbols), when acting on Orlicz modulation spaces. In particular we extend continuity and compactness properties of the Fourier integral operators in [7, 8, 26, 60], where related results were obtained for such operators when acting on classical modulation spaces (of Lebesgue types). These earlier results are in turn extensions and generalizations of pioneering results and investigations performed by Boulkhemair in [5], where $L^{2}$-continuity of subclasses of our Fourier integral operators were established.
@@
is continuous. If, more restrictive, $a$ is chosen in the completion $M^{\sharp,1}_{(\omega)}(\mathbf{R}^{2d+m})$ of $M^{1,1}_{(\omega)}(\mathbf{R}^{2d+m})$ in $M^{\infty,1}_{(\omega)}(\mathbf{R}^{2d+m})$, then the map (0.3) is compact. (See Theorem 2.4.)
We notice that we may choose $\omega$ such that $M^{\infty,1}_{(\omega)}$ contain $S^{r}_{0,0}$. In particular, as announced above, we put less restrictions on the amplitudes compared to the assumptions above adapted to hyperbolic problems.
@@
Secondly, in (0.3) we allow $\omega$, $\omega_{j}$ and $v$ to belong to the large class $\mathscr{P}_{s}$ of weights which are moderate by subexponential functions of degree $s^{-1}<1$, while in [60] it is required that the involed weights should belong to the smaller class $\mathscr{P}$ of polynomially moderate weights. By imposing these two significant restrictions, then our Theorem 2.3 in Section 2, as well as (0.3), essentially takes the form [60, Theorem 2.2].
@@ 1. Preliminaries
In the section we recall some basic facts on Gelfand-Shilov spaces, Orlicz spaces, Orlicz modulation spaces, pseudo-differential operators and Wigner distributions. We also give some examples on Young functions, Orlicz spaces and Orlicz modulation spaces. (See Examples 1.9 and 1.16.) Notice that Young functions are fundamental in the definition of Orlicz spaces and Orlicz modulation spaces).
@@ 1.1. Gelfand-Shilov spaces
By [58, Theorem 2.3] it follows that the definition of the map $(f,\phi)\mapsto V_{\phi}f$ from $\mathscr{S}(\mathbf{R}^{d})\times\mathscr{S}(\mathbf{R}^{d})$ to $\mathscr{S}(\mathbf{R}^{2d})$ is uniquely extendable to a continuous map from $\mathcal{S}_{s}^{\prime}(\mathbf{R}^{d})\times\mathcal{S}_{s}^{\prime}(\mathbf{R}^{d})$ to $\mathcal{S}_{s}^{\prime}(\mathbf{R}^{2d})$, and restricts to a continuous map from $\mathcal{S}_{s}(\mathbf{R}^{d})\times\mathcal{S}_{s}(\mathbf{R}^{d})$ to $\mathcal{S}_{s}(\mathbf{R}^{2d})$. The same conclusion holds with $\Sigma_{s}$ in place of $\mathcal{S}_{s}$, at each occurrence.
In the following proposition we give characterizations of Gelfand-Shilov spaces and their distribution spaces in terms of estimates of the short-time Fourier transform. We omit the proof since the first part follows from [32, Theorem 2.7] and the second part from [58, Theorem 2.5]. See also [15] for related results. Here and in what follows, the notation $A(\theta)\lesssim B(\theta)$, $\theta\in\Omega$, means that there is a constant $c>0$ such that $A(\theta)\leq cB(\theta)$ holds for all $\theta\in\Omega$. We also set $A(\theta)\asymp B(\theta)$ when $A(\theta)\lesssim B(\theta)\lesssim A(\theta)$.
@@ 1.3. Orlicz Spaces
We refer to [62, Lemma 1.18] for the proof of the following lemma.
@@ 1.4. Orlicz modulation spaces
Beside these well-known families of modulation spaces, we shall also consider the modulation space $M^{\sharp,q}_{(\omega)}(\mathbf{R}^{d})$, which consists of all $f\in M^{\infty,q}_{(\omega)}(\mathbf{R}^{d})$ such that
$\lim_{R\to\infty}\left(\left\|\sup_{|x|\geq R}|V_{\phi}f(x,\,\cdot\,)\omega(x,\,\cdot\,)|\right\|_{L^{q}}=0,\right)$
when $q$, $\phi$ and $\omega$ are the same as in Definition 1.10. We notice that $M^{\sharp,1}_{(\omega)}(\mathbf{R}^{d})$ is a central modulation space in [26], and that several invariance properties are deduced in [38]. For example it is here shown that the window function $\phi$ can be any element in a suitable weighted $M^{1,1}$ class. Furthermore, in [38] the following result is obtained.
Lemma 1.11.
Let $\omega\in\mathscr{P}_{E}(\mathbf{R}^{2d})$ and $q\in[1,\infty)$. Then $M^{\sharp,q}_{(\omega)}(\mathbf{R}^{d})$ is the completion of $\Sigma_{1}(\mathbf{R}^{d})$ under the norm $\|\,\cdot\,\|_{M^{\sharp,q}_{(\omega)}}$.
@@
Next we explain some basic properties of Orlicz modulation spaces. The following proposition shows that Orlicz modulation spaces are completely determined by the behavior of the Young functions near origin. We refer to [63, Proposition 5.11] for the proof.
@@ 1.5. Kernel operators and Schatten-von Neumann classes
Throughout the paper, all Hilbert spaces are assumed to be separable, and observe that this is always the case for Hilbert spaces which are continuously embedded in $\mathscr{S}^{\prime}(\mathbf{R}^{d})$, in view of [44, Proposition 1.2]. However, we note that most parts of what is described here also hold when $\mathcal{H}_{1}$ and $\mathcal{H}_{2}$ are allowed to be non-separable.
@@ 2.2. Continuity for Fourier integral operators, when acting on modulation spaces
In order to explain our results on Fourier integral operators, we first give the conditions on involved weight functions. We usually assume that they satisfy
@@
The first result shows that Fourier integral operators with amplitudes in $M^{\infty,1}_{(\omega)}(\mathbf{R}^{2d+N})$ are well-defined as continuous mappings from $\Sigma_{1}(\mathbf{R}^{d})$ to $\Sigma_{1}^{\prime}(\mathbf{R}^{d})$. For polynomial weights, the result is essentially the same as [60, Theorem 2.1]. For general weights, the result follows by similar arguments as for [60, Theorem 2.1]. The proof is therefore omitted.
@@
Let $s>1$, $d_{1}=d_{2}=d$, $\varphi\in C(\mathbf{R}^{2d+m})$, and $\omega,v\in\mathscr{P}_{s}(\mathbf{R}^{2(2d+m)})$ satisfy (0.2), (2.2) and (2.3). Then the map $a\mapsto\operatorname{Op}_{\varphi}(a)$ from $\Sigma_{1}(\mathbf{R}^{2d+m})$ to $\mathcal{L}(\Sigma_{1}(\mathbf{R}^{d}),\Sigma_{1}^{\prime}(\mathbf{R}^{d}))$ extends uniquely to a continuous map from $M^{\infty,1}_{(\omega)}(\mathbf{R}^{2d+m})$ to $\mathcal{L}(\Sigma_{1}(\mathbf{R}^{d}),\Sigma_{1}^{\prime}(\mathbf{R}^{d}))$.
The next two theorems assert continuity and compactness for Fourier integral operators in Theorem 2.2, when they are acting on Orlicz modulation spaces.
Theorem 2.3.
Let $s>1$, $d_{1}=d_{2}=d$, $\Phi$ be a Young function such that $1<q_{\Phi}\leq p_{\Phi}<\infty$, $\omega_{1},\omega_{2}\in\mathscr{P}_{s}(\mathbf{R}^{2d})$, and let $\varphi$, $\omega$, $v$ be the same as in Theorem 2.2. Also let $a\in M^{\infty,1}_{(\omega)}(\mathbf{R}^{2d+m})$. Then $\operatorname{Op}_{\varphi}(a)$ from $\Sigma_{1}(\mathbf{R}^{d})$ to $\Sigma_{1}^{\prime}(\mathbf{R}^{d})$ extends uniquely to a continuous operator from $M^{\Phi}_{(\omega_{1})}(\mathbf{R}^{d})$ to $M^{\Phi}_{(\omega_{2})}(\mathbf{R}^{d})$. Moreover, for some constant $C$ it holds
@@
Theorem 2.4.
Let $s>1$, $d_{1}=d_{2}=d$, $\Phi$ be a Young function such that $1<q_{\Phi}\leq p_{\Phi}<\infty$, $\omega_{1},\omega_{2}\in\mathscr{P}_{s}(\mathbf{R}^{2d})$, and let $\varphi$, $\omega$, $v$ be the same as in Theorem 2.2. Also let $a\in M^{\sharp,1}_{(\omega)}(\mathbf{R}^{2d+m})$. Then $\operatorname{Op}_{\varphi}(a)$ from $M^{\Phi}_{(\omega_{1})}(\mathbf{R}^{d})$ to $M^{\Phi}_{(\omega_{2})}(\mathbf{R}^{d})$ is compact.
Theorem 2.4 will follow by combining Theorem 2.3, with certain density arguments and the following proposition.
Proposition 2.5.
Let $s>1$, $\Phi$, $\varphi$, $\omega$, $v$, $\omega_{1}$, $\omega_{2}$ be the same as in Theorem 2.3, and let $a\in\Sigma_{1}(\mathbf{R}^{2d+m})$. Then $\operatorname{Op}_{\varphi}(a)$ from $M^{\Phi}_{(\omega_{1})}(\mathbf{R}^{d})$ to $M^{\Phi}_{(\omega_{2})}(\mathbf{R}^{d})$ is compact.
Before proving Theorems 2.3, 2.4, and Proposition 2.5, we present some further continuity properties for Fourier integral operators. The amplitudes should fulfill norm estimates, with norm $\|\,\cdot\,\|$ either defined by
@@
Then [60, Lemma 2.2] shows that
@@
We shall follow the proof of [60, Theorem 2.1] when proving (2). If $\Phi(t_{0})=0$ for some $t_{0}>0$, then $M^{\Phi}_{(\omega_{j})}(\mathbf{R}^{d})=M^{\infty}_{(\omega_{j})}(\mathbf{R}^{d})$, and the result follows from [60, Theorem 2.1]. Therefore suppose that $\Phi(t)>0$ when $t>0$. Then $\mathscr{S}(\mathbf{R}^{d})$ is dense in $M^{\Phi}_{(\omega_{j})}(\mathbf{R}^{d})$.
@@
Proof of Proposition 2.5.
Let $X=(x,y,\zeta)$,
$\vartheta(X,\xi,\eta,z)$ $=\omega(X,\xi,\eta,z)(1+|x|+|y|+|z|+|\xi|)^{3}$
and
$\vartheta_{2}(x,\xi)$ $=\omega_{2}(x,\xi)(1+|x|)(1+|\xi|).$
Since $|\varphi^{\prime\prime}(x,y,\zeta)|$ is a bounded function, it follows that
$|\varphi^{\prime}_{x}|=|\varphi^{\prime}_{x}(X)|\lesssim(1+|x|+|y|+|\zeta|).$
This gives
$\omega(X,\xi-\varphi^{\prime}_{x},\eta-\varphi^{\prime}_{y},-\varphi^{\prime}_{\zeta})(1+|x|)(1+|\xi|)$
$\leq\omega(X,\xi-\varphi^{\prime}_{x},\eta-\varphi^{\prime}_{y},-\varphi^{\prime}_{\zeta})(1+|x|)(1+|\xi-\varphi^{\prime}_{x}|)(1+|\varphi^{\prime}_{x}|)$
$\lesssim\omega(X,\xi-\varphi^{\prime}_{x},\eta-\varphi^{\prime}_{y},-\varphi^{\prime}_{\zeta})(1+|x|+|y|+|\zeta|+|\xi-\varphi^{\prime}_{x}|)^{3}$
$=\vartheta(X,\xi-\varphi^{\prime}_{x},\eta-\varphi^{\prime}_{y},-\varphi^{\prime}_{\zeta}).$
A combination of these estimates and (2.2) shows that (2.2) holds with $\vartheta$ and $\vartheta_{2}$ in place of $\vartheta$ and $\vartheta_{2}$.
Since $a\in\Sigma_{1}(\mathbf{R}^{2d+N})$, it follows that $a\in M^{\infty,1}_{(\vartheta)}(\mathbf{R}^{2d+N})$. Hence Theorem 2.3 shows that
$\operatorname{Op}_{\varphi}(a):M^{\Phi}_{(\omega_{1})}(\mathbf{R}^{d})\to M^{\Phi}_{(\vartheta_{2})}(\mathbf{R}^{d})$ (2.22)
is continuous.
Since the inclusion map
$\boldsymbol{\iota}:M^{\Phi}_{(\vartheta_{2})}(\mathbf{R}^{d})\to M^{\Phi}_{(\omega_{2})}(\mathbf{R}^{d}),$ (2.23)
is compact, in view of [61, Proposition 4.8], it follows that
$\operatorname{Op}_{\varphi}(a)=\boldsymbol{\iota}\circ\operatorname{Op}_{\varphi}(a):M^{\Phi}_{(\omega_{1})}(\mathbf{R}^{d})\to M^{\Phi}_{(\omega_{2})}(\mathbf{R}^{d})$
is compact, giving the result.
@@
By Theorem 2.3 we have
$\|\operatorname{Op}_{\varphi}(a)\|_{M^{\Phi}_{(\omega_{1})}\to M^{\Phi}_{(\omega_{2})}}\lesssim\|a\|_{M^{\infty,1}_{(\omega)}}.$
Since $a\in M^{\sharp,1}_{(\omega)}(\mathbf{R}^{2d+N})$, it follows from Lemma 1.11 and Proposition 2.5, that for some $a_{j}\in\Sigma_{1}(\mathbf{R}^{2d+N})$, one has
$\|a-a_{j}\|_{M^{\infty,1}_{(\omega)}}\to 0,\quad\text{as}\quad j\to\infty.$
A combination of these estimates gives
$\|\operatorname{Op}_{\varphi}(a)-\operatorname{Op}_{\varphi}(a)\|_{M^{\Phi}_{(\omega_{1})}\to M^{\Phi}_{(\omega_{2})}}\lesssim\|a-a_{j}\|_{M^{\infty,1}_{(\omega)}}\to 0,\quad\text{as}\quad j\to\infty.$
Hence, $\operatorname{Op}_{\varphi}(a)$ can be approximated in norm by $\operatorname{Op}_{\varphi}(a_{j})$. Since $a_{j}\in\Sigma_{1}(\mathbf{R}^{2d+N})$, it follows that $\operatorname{Op}_{\varphi}(a_{j})$. A combination of these properties shows that $\operatorname{Op}_{\varphi}(a)$ is compact, and the result follows. ∎
Proof of Theorem 2.6.
The result follows from [60, Theorem 2.7] when $p_{\Phi}=q_{\Phi}$. Therefore we assume that $1<q_{\Phi}\leq p_{\Phi}<\infty$.